sues mother is 22 years older than susan and the sum of their ages is 60. find their ages
Answer by lenny460(1073)

sues mother is 22 years older than susan and the sum of their ages is 60. find their ages

Let the ages be:
Susan = n
Susan's mother = n + 22
n + n + 22 = 60
2n + 22 = 60
Subtract 22 from each side of the equation.
2n + 22 - 22 = 60 - 22
2n = 38
n = 19
n + 22 = 19 + 22 = 41
Therefore:
Susan is 19 years old.
Susan's mother is 41 years old.




Lennox Obuong
Algebra Tutor